Helen and James had 160 CDs between them Helen said to James “if I had five more CDs I would have twice as many as you” how many CDs do they each have?
Answer by ikleyn(52787)

.
H + 5 = 2*(160-H)


H + 5 = 320 - 2H


H + 2H = 320 - 5


3H = 315


H = 315/3 = 105.


ANSWER.  Helen has 105 CDs.  James has the rest, 160 - 105 = 55 CDs.